1 / 7

Beam Optimization and Performance Update for LHC Fill #2628 on May 14th, 2023

On May 14th, 2023, significant developments occurred during the LHC operations related to beam performance and optimization. The day included stabilization of beams with a peak luminosity reaching 4.8e33 cm-2s-1, along with a notable re-steering of the beam and various checks on the alignment of key components. Issues with beam quality were also investigated, particularly during the transfer from PS to SPS. Further efforts included adjustments to beam scraping techniques to minimize losses and ensure physical integrity. Continuous monitoring of beam positions revealed drifting, necessitating corrective measures.

nerita
Télécharger la présentation

Beam Optimization and Performance Update for LHC Fill #2628 on May 14th, 2023

An Image/Link below is provided (as is) to download presentation Download Policy: Content on the Website is provided to you AS IS for your information and personal use and may not be sold / licensed / shared on other websites without getting consent from its author. Content is provided to you AS IS for your information and personal use only. Download presentation by click this link. While downloading, if for some reason you are not able to download a presentation, the publisher may have deleted the file from their server. During download, if you can't get a presentation, the file might be deleted by the publisher.

E N D

Presentation Transcript


  1. Monday 14th May • 00:03UFO at BLMQI.06R8.B2E10_MQML. Beam dumpRe-steering required for TI2 • 03:28 STABLE BEAMS #2627. 1.27e11 ppb, 4.11e33 cm-2s-1 • 13:36 Operator dump fill #2627. Physics for 10:08 hours. Intlumi100 pb-1 • 15:19 TDI measurement - check alignment of jaws 'warm' • 16:30 Optimising beam in injectors…. • 20:55 Injecting for physics, fill #2628 • 22:17 Stable beams, initial lumi’s 4.8e33 cm-2s-1 • 06:07 Beams dumped, trip of ROF.A67B1 • Intlumi 97 pb-1 • Access for power converter piquet LHC 8:30 meeting

  2. Fill #2627 OK next fill, o-night • Dumped by operator after 10:08 hours and about 100 pb-1 • Around 8:00 start of • IP8 beam positions seem to be drifting • Drifts of luminous region position observed at IP1,5,8 (LPC) • Sharp increase of lumi in IP2 -> increase of satellites • Occasional ‘noisy RF’ requiring some gentle abort gap cleaning • Philippe investigating LHC 8:30 meeting

  3. Check TDI alignment ‘warm’ Tuesday last weekInner limit (RU) increased by 50 um Measurement IL Measured LVDT RU position cold • No interlocks on TDI IP8 position anymore since last week • The RU LVDT interlock was increased by 50 um • During physics at 55 mm instead of 37.5 mm: less heating • Want to check alignment with beam • Lever arm effect, don’t know the real position of jaw & jumps position. • Found jaw positions for IP2 and IP8 between 6.4 – 6.6 sigma. • Setting 6.8: OK LHC 8:30 meeting

  4. Fill #2628 Federico Roncarolo • Problems with beam quality in transfer from PS – SPS • Suspect extraction in PS • Could not really fix, investigations will need to continue • SPS also ‘optimised’, likely solved the difference between batches • Beam sizes at injection with BSRT in test mode: flatter LHC 8:30 meeting

  5. Lumi per crossing also flatter From CMS LHC 8:30 meeting

  6. Fill #2628 During ramp TCSG.A6R7.B2, integration time: 83.8 s, losses of 34 % of dump limit Slightly less losses then in previous fill Higher lumi’s: around 4.8e33 cm-2s-1 Losses during physics up to 40 % of dump threshold, IP5 and IP8 LHC 8:30 meeting

  7. Plan / Outstanding • Plan • Finish access for PO, ROFA67.B1 – until at least 9:30 & pre-cycle • Scraping at injection to reduce losses in squeeze • Physics • Outstanding • Continue investigations beam quality from injector chain • Orbit in collision needs to be corrected and trims incorporated (in particular point 5) • Possibly LHCb polarity change on Wednesday’ LHC 8:30 meeting

More Related